An inclined plane of inclination (&) has a length of 15m .If a body rolls down the plane freely along the entire length with an acceleration 5ms^-2 , calculate the
(a)angle (&)
(b)height of the plane
NB:& = theta

Take g=10m/s²
1
Expert's answer
2020-07-02T17:13:41-0400

a)


"a=g\sin{\theta}\\5=10\sin{\theta}\\\theta=30\degree"

b)


"f=15\sin{30}=7.5\ m"